OpenStudy (elkiller33):

write 7/50 as a decimal

OpenStudy (naate):

.14

OpenStudy (naate):

ohh wait sorry

OpenStudy (ebonygotthejuice):

= 7/50 = 7 ÷ 50 = 0.14

OpenStudy (naate):

ok, so since its 7/50, what do you have to do to get 50 to equal 100?

OpenStudy (karen.joyce.bittner):

.14 just divide 7/50 and you get .14 hope that helped